Presentation on theme: "EMCal DCS Status By B. S. Nilsen, M. Cherney, J. Fujita Creighton University For the EMCal project."— Presentation transcript:
EMCal DCS Status By B. S. Nilsen, M. Cherney, J. Fujita Creighton University For the EMCal project
Contents Problems to be solved Completed tasks Plans
Problems to be Solved Introduction of Recipes –Been having difficulties making a working implementation. –Recent workshop been very helpful. –Want to have a backup dcs system before we try implementing this again. –Needed to allow for multiple configurations like LED running. Busy occurring during Physics run attempts –Not seen during stand alone runs –Effects ERP Drill and LHC Dry runs ERP Drill and LHC Dry run –Most recent attempt when well (ERP), but effected by busy
Completed Tasks 4/12 Super-modules installed and working. Emergency Button implemented. L0 timing issue understood and patched.
Plans Continue to tack down Busy problems. Obtain and install Parallel PVSS system on new DCS machine. –Implement recipes. –Improve and streamline Control system. –Set up Cooling control and monitoring. –Implement LED calibration. Reestablish remote development and test bed.
The End Thats All Folks
Outline Installation Fee Cards DCS machines Recipes Cooling Emergency Button PVSS panels Going Busy Physics Drill High voltage –ISEG OPC/CANHVcontrol Low voltage Led calibration system L0 timing Planning LHC Dry Run
Installation 4 out of 12 Super modules installed and working.
L0 Latency L0 presently arriving 960ns after interaction (352 BC) Expect to improve by 6 or 7 BC (150- 175 ns)
Emergency Button Installed in August Hardware version yet to come.
Fee cards Suspected noise fee card has been swapped (twice now) but noise still there. Problem must be else where. –LTU system. –ELMB system. –?
Going Busy PVSS has Intermittently been going Busy, when trying to go global, OK when in standalone run. Is SOD checks in the physics runs? Still under investigation.
Recipes Been having difficulties making a working implementation. Recent workshop been very helpful. Want to have a backup dcs system before we try implementing this again. Needed to allow for multiple configurations like LED running
Cooling Still some hardware problems and developments under way. PVSS monitoring still under development.
DCS Machine Present machine limits being reached. Acquisition of additional machine just starting. CERN has new contract with HP and dcs machine parameters still under negotiations.
High voltage Intermittent Problems with Iseg HV system have been seen. Simultaneous running of Iseg OPC and IsegCANHVcontrol may be source of problem. Fix planned ASAP.
LHC Dry Run First Try: Failed due to problems with LHC clock switching 10/16/09 Was busy and not in today. 10/19/09 Central DCS has all Detectors In READY state All 4 LHC Experiments will perform 2*(READY-BEAMTUNING-READY)
Physics (ERP) Drill First attempt took 3 tries 10/15/09 Attempt EMCal failed. Most resent drill Performed OK Cycle through READY BEAM-Tuning READY
LED System LED are now working Will require changes in PVSS for proper operations (recipes).
Low Voltage System Operational
PVSS Panels System to date is adequate Will need to be redesigned to better support of full EMCal. Will attempt to make operational efficiencies.
Planning Iseg OPC CANHVcontrol conflict to be removed (go to all OPC control). Continue to tack down Busy problems. Obtain and install Parallel PVSS system on new DCS machine. –Implement recipes. –Improve and streamline Control system. –Set up Cooling control and monitoring. –Implement LED calibration. Reestablish remote development and test bed.